# Pre-training Cross-lingual Open Domain Question Answering with Large-scale Synthetic Supervision

### v1

@@
Cross-lingual question answering (CLQA) is a complex problem, comprising cross-lingual retrieval from a multilingual knowledge base, followed by answer generation either in English or the query language. Both steps are usually tackled by separate models, requiring substantial annotated datasets, and typically auxiliary resources, like machine translation systems to bridge between languages. In this paper, we show that CLQA can be addressed using a single encoder-decoder model. To effectively train this model, we propose a self-supervised method based on exploiting the cross-lingual link structure within Wikipedia. We demonstrate how linked Wikipedia pages can be used to synthesise supervisory signals for cross-lingual retrieval, through a form of cloze query, and generate more natural queries to supervise answer generation. Together, we show our approach, CLASS, outperforms comparable methods on both supervised and zero-shot language adaptation settings, including those using machine translation.
@@ 1 Introduction
Рис. 1: The overview of our two-stage unsupervised pre-training method for cross-lingual open domain question answering. Note that translations in stage 2 are employed for illustrative purposes only and are not used for training.
@@
This large-scale pre-training framework empowers the model to demonstrate promising unsupervised performance, and it can even outperform many competitive supervised counterparts. By fine-tuning it with supervised English and multilingual QA data, we can attain further improvements, ultimately establishing new state-of-the-art performance in both cross-lingual retrieval and multilingual open-domain QA tasks. In summary, our contributions are: [^1] Source code will be released at https://github.com/Fantabulous-J/CLASS
@@ 2.1 Task Definition
Given a question $q^{L}$ in language $L$, Cross-lingual Passage Retrieval requires retrieving a collection of passages $\mathcal{D}^{En}$ from English Wikipedia $\bm{C^{En}}$ that potentially provide evidence to answer $q^{L}$. An English answer $a^{En}$ is then generated by a cross-lingual reader with retrieved information. In contrast, Multilingual Open-Domain Question Answering aims at answering $q^{L}$ in language $L$ by referring to a multilingual Wikipedia $\bm{C^{Multi}}$. In this setting, the prior knowledge of which language contains the evidence is unavailable, and the relevant passages can be retrieved from any language.
@@ 2.2 Model Architecture
We use a single encoder-decoder language model to perform both passage retrieval and answer generation task (Lee et al., 2022; Jiang et al., 2022).
Retriever
We divide the encoder into two parts, with the first $B$ layers used as the dual-encoder, where the query and passage are encoded independently. Following Jiang et al. (2022), we use the query $Q$ and key vectors $K$ in $B+1$-th layer as query and passage representations, respectively:
$E_{\mathbf{d}}=\{K_{\mathbf{d}}^{B+1,h}\in\mathbb{R}^{|\mathbf{d}|\times e}\}_{h=1}^{H},$
$E_{\mathbf{q}}=\{Q_{\mathbf{q}}^{B+1,h}\in\mathbb{R}^{|\mathbf{q}|\times e}\}_{h=1}^{H},$
where $|\mathbf{q}|$ and $|\mathbf{d}|$ are query and passage length. $H$ is the number of self-attention heads and $e$ is the dimension of each head.
The self-attention matrix $\operatorname{SA}_{\mathbf{q},\mathbf{d}}^{B+1,h}$ from a specific head is considered the source of retrieval scores. A sum of max computations (Khattab and Zaharia, 2020) are performed to reduce it to yield the retrieval score:
$s(q,d)$ $=\sum_{i\in|\mathbf{q}|}\max_{j\in|\mathbf{d}|}\operatorname{SA}_{i,j}^{B+1,h},$
$\operatorname{SA}_{\mathbf{q},\mathbf{d}}^{B+1,h}$ $=Q_{\mathbf{q}}^{B+1,h}\times K_{\mathbf{d}}^{B+1,h^{\top}}\in\mathbb{R}^{|\mathbf{q}|\times|\mathbf{d}|}.$
Reader
After the retriever layer, both query and passage are encoded jointly, wherein the remaining encoder layers function as the cross-encoder. Then, the joint encodings of multiple passages $\{E_{\mathbf{q,d_{i}}}\}_{i=0}^{n}$ are integrated into the decoder to generate the answer efficiently following the Fusion-in-Decoder approach (Izacard and Grave, 2021).
2.3 Training Objectives
The model is trained end-to-end by employing the training signals derived from the answer generation task: $\mathcal{L}_{\text{e2e}}=\mathcal{L}_{\text{reader}}+\alpha\cdot\mathcal{L}_{\text{KL}}$.
Reader Loss
The reader is trained with language modelling loss via teacher forcing, by providing the query $q$ and a collection of evidence $\mathcal{D}$ as input:
$\mathcal{L}_{\text{reader}}=P_{\text{ans}}(a|q,\mathcal{D})=-\log\prod_{t=1}^{T}P(a_{t}|a_{<t},q,\mathcal{D})$
Retriever Loss
The retriever is trained by learning from the decoder, wherein the cross-attention score serves as the passage relevance measurement:
$\mathcal{L}_{\text{KL}}$ $=\mathbb{KL}(P_{\text{be}}(\cdot|q,\mathcal{D})||P_{\text{ca}}(\cdot|q,\mathcal{D})),$
$P_{\text{be}}(\cdot|q,\mathcal{D})$ $=\frac{\exp(s(q,d_{i}))}{\sum_{d_{j}\in\mathcal{D}}\exp(s(q,d_{j}))}\ |\ d_{i}\in\mathcal{D},$
$P_{\text{ca}}(\cdot|q,\mathcal{D})$ $=\sum_{h=0}^{H}\sum_{t=0}^{|d_{i}|}\frac{\operatorname{SG}(\operatorname{CA}(0,h,t))}{H}\ |\ d_{i}\in\mathcal{D},$
where $P_{\text{be}}(\cdot|q,\mathcal{D})$ and $P_{\text{ca}}(\cdot|q,\mathcal{D})$ are the distributions over $\mathcal{D}$ from the retriever and the decoder’s cross-attention scores, respectively. $\operatorname{SG}$ signifies the stop-gradient operation, and $\operatorname{CA}$ denotes the cross-attention score at the last decoder layer. The term 0 refers to the first output token, $H$ is the number of cross-attention heads, and $|d_{i}|$ stands for the length of passage $d_{i}$.
@@ 3 Method
We propose an unsupervised two-stage pre-training method for cross-lingual open-retrieval question answering, as depicted in Figure 1. Our approach starts with cross-lingual retrieval pre-training, where the multilingual model develops excellent cross-lingual dense retrieval capabilities. This proficiency is acquired through learning from a well-trained English model, employing cloze-style parallel queries and retrieved English passages as inputs. The subsequent stage involves pre-training for multilingual question-answering (QA), where the model is further pre-trained on multilingual question-answer pairs that are automatically generated. This process entails selecting potential answers from anchor texts and applying our novel question transformation techniques to convert cloze questions into natural questions by prompting a large language model.
@@ 3.1 Cross-Lingual Retrieval Pre-training
$\operatorname{M}(x,y)=\frac{cos(x,y)}{\sum_{z\in N_{x}}\frac{cos(x,z)}{2k}+\sum_{z\in N_{y}}\frac{cos(y,z)}{2k}},$
where $N_{x}$ and $N_{y}$ are the top-$k$ neighbours of sentence $x$ and $y$ in the other language, respectively. $cos(x,y)$ denotes the cosine similarity between the embeddings of $x$ and $y$. We follow mSimCSE (Wang et al., 2022) to pre-train a multilingual model with unsupervised contrast learning and employ it for embedding extraction. We apply this scoring function to $q_{i}^{En}$ and each $q_{j}^{L}\in\{q_{j}^{L}\}_{j=0}^{n}$. Pairs whose scores surpass a pre-defined threshold $T$ are selected as parallel queries, denoted as $\{q_{i}^{En},q_{j}^{L},a_{i}\}$. [^2] We identify $a_{i}$ and mask it in $q_{j}^{L}$ through string match if $L$ is written in Latin script and leave $q_{j}^{L}$ unchanged otherwise.
Training
A well-trained English model $\theta^{En}$ is employed to teach a multilingual model $\theta^{ML}$ using parallel queries. Specifically, given a training example $\{q^{En},q^{L},a\}$, we employ $\theta^{En}$ to retrieve a set of relevant passages $\mathcal{D}^{En}$ from English Wikipedia $\bm{C^{En}}$ for $q^{En}$. The multilingual model is then compelled to align its retrieval distributions with those of $\theta^{En}$ over $\mathcal{D}^{En}$ through KL divergence loss:
@@
$=\mathbb{KL}(P^{ML}_{\text{be}}(\cdot|q^{L},\mathcal{D}^{En})||P^{En}_{\text{be}}(\cdot|q^{En},\mathcal{D}^{En}))$
@@
$\mathcal{L}_{\text{reader}}=P_{\text{ans}}(a|q^{En},\mathcal{D}^{En})+P_{\text{ans}}(a|q^{L},\mathcal{D}^{En}).$
@@
$=\mathbb{KL}(P^{ML}_{\text{be}}(\cdot|q^{L},\mathcal{D}^{En})||P^{ML}_{\text{be}}(\cdot|q^{En},\mathcal{D}^{En}))$
@@ 3.2 Multilingual QA Pre-training
The cloze questions are substantially different from the formats of natural questions, which inherently impedes the development of advanced question-answering skills. Moreover, the incapacity to precisely locate and mask the answer $a$ within $q^{L}$ for perfectly aligned queries makes the QA task notably simpler, as $a$ implicitly appears in $q^{L}$ (e.g., "インド"in $q^{L}$ is the Japanese answer in Figure 1). Meanwhile, since $q^{En}$ and $q^{L}$ could be roughly aligned, the querying of $a$ by $q^{L}$ is not assured, thereby introducing noise into the pre-trained data (e.g., "In 1945, his father sent him to Collège des Frères"and "父はサブリーをヤッファのカトリック系フランス語学校に送った。"are aligned but the Japanese query does not mention the answer 1945). Thus, we design another pre-training technique to address the limitations above.
@@ 3.2.1 Pre-training Data
Initial Data
In contrast to English texts, where robust NER systems facilitate the detection of named entities with high precision for answer generation, such systems in other languages exhibit inherent deficiencies. Instead, we employ anchor texts with hyperlinks as answer candidates. Specifically, for a given sentence $s^{L}$ in language $L$, we consider the anchor texts $\{a_{i}^{L}\}_{i=0}^{n}$ within it as potential answers and construct cloze questions $\{q_{i}^{L}\}_{i=0}^{n}$ accordingly.
For each $a_{i}^{L}$, we fetch the Wikipedia page $\mathcal{W}^{L}$ to which it links and access the corresponding English Wikipedia page $\mathcal{W}^{En}$ via language link. Subsequently, the title $a_{i}^{En}$ of $\mathcal{W}^{En}$ is assumed to be the pseudo translation of $a_{i}^{L}$. Moreover, NER tagging is performed on the first paragraph of $\mathcal{W}^{En}$ to identify the type $t_{i}$ of the title entity $a_{i}^{En}$, which is then assigned to $a_{i}^{L}$. Finally, a training example is derived in the form of $(q_{i}^{L},a_{i}^{L},a_{i}^{En},t_{i})$. [^3] Notably, $a_{i}^{L}$ is chosen for training when the downstream task involves multilingual question-answering. Likewise, $a_{i}^{En}$ is employed when English answer generation is of interest.
Query Transformation
@@
where wh_word is chosen according to the entity type $t_{i}$ through heuristics (Lewis et al., 2019). This step yields a curated set of ICL examples: $\mathbb{K}=\{q_{i}^{L},\texttt{wh\_word},a_{i}^{L},{q_{i}^{L}}^{T}\}_{i=0}^{k}$. An example is shown in Figure 9 in the Appendix.
We employ an open-sourced LLM, LLaMA-2-7B (Touvron et al., 2023), for the efficient transformation of questions at large scale. Specifically, the prompt is constructed as follows: (1) the first part includes the instruction that expresses the intent of the question transformation task; (2) $n$ instances are randomly sampled from $\mathbb{K}$ and organised in a pre-defined structure; (3) the last part includes the test instance to be transformed. Examples are shown in Appendix D.
3.2.2 Training
For each $({q^{L}}^{T},a^{L})$, we use the retrieval component of the current model $\theta^{ML}$ to gather a set of passages $\mathcal{D}^{Multi}$ from $\bm{C^{Multi}}$, [^4] We replace $a^{L}$ with $a_{i}^{En}$ and $\bm{C^{Multi}}$ with $\bm{C^{En}}$ when the downstream task of interest is English answer generation. and optimise the model using $\mathcal{L}_{\text{e2e}}$ as defined in §2.3. To ensure efficient training, passages are retrieved for each training query beforehand, with periodic refreshing at specific intervals using the most recent model.

### v3

@@
Cross-lingual open domain question answering (CLQA) is a complex problem, comprising cross-lingual retrieval from a multilingual knowledge base, followed by answer generation in the query language. Both steps are usually tackled by separate models, requiring substantial annotated datasets, and typically auxiliary resources, like machine translation systems to bridge between languages. In this paper, we show that CLQA can be addressed using a single encoder-decoder model. To effectively train this model, we propose a self-supervised method based on exploiting the cross-lingual link structure within Wikipedia. We demonstrate how linked Wikipedia pages can be used to synthesise supervisory signals for cross-lingual retrieval, through a form of cloze query, and generate more natural questions to supervise answer generation. Together, we show our approach, CLASS, outperforms comparable methods on both supervised and zero-shot language adaptation settings, including those using machine translation.
@@ 1 Introduction
Рис. 1: The overview of our two-stage unsupervised pre-training method for cross-lingual open domain question answering. English translations from Google Translate are added in (b) for readability.
@@
This large-scale pre-training framework empowers the model to demonstrate promising unsupervised performance, and it can even outperform many competitive supervised counterparts. By fine-tuning it with supervised English and multilingual QA data, we can attain further improvements, ultimately establishing new state-of-the-art performance in both cross-lingual retrieval and multilingual open-domain QA tasks. In summary, our contributions are: [^1] Code and data are available here.
@@ 2.1 Task Definition
Given a query $q^{L}$ in language $L$, Cross-lingual Passage Retrieval requires retrieving a collection of passages $\mathcal{D}^{En}$ from English Wikipedia $\bm{C^{\textbf{En}}}$ that potentially provide evidence to answer $q^{L}$. In contrast, Multilingual Open-Domain Question Answering aims at answering $q^{L}$ in language $L$ by referring to a multilingual Wikipedia $\bm{C^{\textbf{Multi}}}$. In this setting, the prior knowledge of which language contains the evidence is unavailable, and the relevant passages can be retrieved from any language.
Рис. 2: The unified model for passage retrieval and question answering.
@@ 2.2 Model Architecture
Figure 2 shows the overall structure of our model. In this model, the bottom layers of the encoder function as the retriever, encoding queries and passages independently for efficient retrieval. The remaining encoder layers and the entire decoder are designated as the reader for question answering.
Retriever.
The retriever is a bi-encoder that uses the first $B$ encoder layers with $H$ heads to encode query $q$ and passages $d$ from a corpus $\mathcal{D}$. We use the query $Q$ and key vectors $K$ in $B+1$-th layer as their embeddings, respectively (Jiang et al., 2022):
where $|\mathbf{q}|$ and $|\mathbf{d}|$ are sequences lengths and $e$ is the dimension of each head.
The self-attention matrix $\operatorname{SA}_{\mathbf{q},\mathbf{d}}^{B+1,h}$ from a specific head ($h=6$ (Jiang et al., 2022)) is considered the source of retrieval scores. A sum of max computations (Khattab and Zaharia, 2020) is performed to reduce it to yield the retrieval score:
We denote this as Multi-Vector Retrieval and consider it as our default setting. We also explore Dense Retrieval, which takes the average pooling of layer $B$’s output with $\operatorname{LayerNorm}$ as query $Q_{\mathbf{q}}$ and passage $K_{\mathbf{d}}$ representations, and the relevance is measured by their dot product:
$s_{\text{dense}}(q,d)=\operatorname{LN}(Q_{\mathbf{q}})\cdot\operatorname{LN}(K_{\mathbf{d}}).$
The top-$k$ most relevant passages are then retrieved by $\mathcal{D}_{q}=\arg\operatorname{topk}_{d_{i}\in\mathcal{D}}P_{\text{be}}(\cdot|q,D)=\arg\operatorname{topk}\left[s(q,d_{0}),\dots,s(q,d_{|{d\in\mathcal{D}}|})\right]$.
Reader.
The encoded query and each top-$k$ passage in $\mathcal{D}_{q}$ are concatenated and fed into the remaining cross-encoder layers. Finally, the joint encodings $\{E_{\mathbf{q,d_{i}}}\}_{i=0}^{|\mathcal{D}_{q}|}$ are integrated into the decoder through cross-attention to generate the answer $a$ efficiently (Izacard and Grave, 2021b): $P_{\text{ans}}(a|q,\mathcal{D}_{q})=\log\prod_{t=1}^{T}P(a_{t}|a_{<t},q,\mathcal{D}_{q})$.
@@ 3 Method
We propose an unsupervised two-stage pre-training method for cross-lingual open-retrieval question answering, as depicted in Figure 1. Our approach starts with cross-lingual retrieval pre-training, where the unified multilingual model develops excellent cross-lingual dense retrieval capabilities. This proficiency is acquired through learning from a well-trained English model, employing cloze-style parallel queries and retrieved English passages as inputs. The subsequent stage involves pre-training for multilingual question-answering (QA), where the unified model is further pre-trained on multilingual question-answer pairs that are automatically generated. This process entails selecting potential answers from anchor texts and applying our novel question transformation techniques to convert cloze questions into natural questions by prompting a large language model.
@@ 3.1 Cross-Lingual Retrieval Pre-training
$\operatorname{M}(q_{i},q_{j})=\frac{\cos(q_{i},q_{j})}{\sum_{z\in N_{q_{i}}}\frac{\cos(q_{i},z)}{2k}+\sum_{z\in N_{q_{j}}}\frac{\cos(q_{j},z)}{2k}},$
where $N_{q_{i}}$ and $N_{q_{j}}$ are the top-$k$ neighbours of sentence $q_{i}$ and $q_{j}$ in the other language, respectively. $cos(q_{i},q_{j})$ denotes the cosine similarity between the embeddings of $q_{i}$ and $q_{j}$ extracted using mSimCSE (Wang et al., 2022). We apply this scoring function to $q_{i}^{En}$ and each $q_{j}^{L}\in\{q_{j}^{L}\}_{j=0}^{n}$. Pairs whose scores surpass a threshold $T$ are selected as parallel queries, denoted as $\{q_{i}^{En},q_{j}^{L},a_{i}\}$. [^2] We identify $a_{i}$ and mask it in $q_{j}^{L}$ through string match if $L$ is written in Latin script and leave $q_{j}^{L}$ unchanged otherwise.
Training.
A well-trained English model $\theta^{En}$ is employed to teach a multilingual model $\theta^{ML}$ using parallel queries. Specifically, given a training example $\{q^{En},q^{L},a\}$, we employ $\theta^{En}$ to retrieve a set of relevant passages $\mathcal{D}_{q^{En}}$ from English Wikipedia $\bm{C^{\textbf{En}}}$ for $q^{En}$. The multilingual model is then compelled to align its retrieval distributions with those of $\theta^{En}$ over $\mathcal{D}_{q^{En}}$ through KL divergence loss:
@@
$=\mathbb{KL}(P^{ML}_{\text{be}}(\cdot|q^{L},\mathcal{D}_{q^{En}})||P^{En}_{\text{be}}(\cdot|q^{En},\mathcal{D}_{q^{En}}))$
@@
$\mathcal{L}_{\text{reader}}=-P_{\text{ans}}(a|q^{En},\mathcal{D}_{q^{En}})-P_{\text{ans}}(a|q^{L},\mathcal{D}_{q^{En}}).$
@@
$=\mathbb{KL}(P^{ML}_{\text{be}}(\cdot|q^{L},\mathcal{D}_{q^{En}})||P^{ML}_{\text{be}}(\cdot|q^{En},\mathcal{D}_{q^{En}}))$
@@ 3.2 Multilingual QA Pre-training
The cloze questions used in §3.1 are substantially different from the formats of natural questions asked by real users, which inherently impedes the development of advanced QA skills. Moreover, the incapacity to precisely locate and mask the answer $a$ within $q^{L}$ for perfectly aligned queries makes the QA task notably simpler, as $a$ implicitly appears in $q^{L}$ (e.g., "{CJK}UTF8minインド"in $q^{L}$ is the Japanese answer in Figure 1 (a)). Meanwhile, since $q^{En}$ and $q^{L}$ could be roughly aligned, the querying of $a$ by $q^{L}$ is not assured, thereby introducing noise into the pre-trained data (e.g., "In 1945, his father sent him to Collège des Frères"and "{CJK}UTF8min父はサブリーをヤッファのカトリック系フランス語学校に送った。"are aligned but the Japanese query does not mention the answer 1945). Thus, we design another pre-training technique to address the limitations above.
@@ 3.2.1 Pre-training Data
Initial Data.
In contrast to English texts, where robust NER systems facilitate the detection of named entities with high precision for answer generation, such systems in other languages exhibit inherent deficiencies. Instead, we employ anchor texts with hyperlinks as answer candidates. Specifically, for a given sentence $s^{L}$ in language $L$, we consider the anchor texts $\{a_{i}^{L}\}_{i=0}^{n}$ within it as potential answers and construct cloze questions $\{s_{i}^{L}\}_{i=0}^{n}$ accordingly.
For each $a_{i}^{L}$, we fetch the Wikipedia page $\mathcal{W}^{L}$ to which it links and access the corresponding English Wikipedia page $\mathcal{W}^{En}$ via language link. Subsequently, the title $a_{i}^{En}$ of $\mathcal{W}^{En}$ is assumed to be the pseudo translation of $a_{i}^{L}$ (Figure 1 (b)). Moreover, NER tagging is performed on the first paragraph of $\mathcal{W}^{En}$ to identify the type $t_{i}$ of the title entity $a_{i}^{En}$, which is then assigned to $a_{i}^{L}$. Finally, a training example is derived as $(s_{i}^{L},a_{i}^{L},a_{i}^{En},t_{i})$.
Query Transformation.
@@
where wh_word is chosen according to the entity type $t_{i}$ through heuristics (Lewis et al., 2019). This step yields a curated set of ICL examples: $\mathbb{K}=\{c_{i}^{L},\texttt{wh\_word},a_{i}^{L},q_{i}^{L}\}_{i=0}^{k}$. An example is shown in Figure 11 in the Appendix.
Subsequently, the curated ChatGPT examples are used as the source to few-shot prompt a smaller LLM, LLaMA-2-7B (Touvron et al., 2023), to generate many more instances efficiently. We include the prompting examples in Appendix E.
3.2.2 Joint Training
The retriever learns indirectly from the answer generation task, taking the cross-attention score from the decoder as the target for query-passage relevance measurement (Izacard and Grave, 2021a):
$\mathcal{L}_{\text{KL}}=\mathbb{KL}(P_{\text{be}}(\cdot|q^{L},\mathcal{D}_{q^{L}})||P_{\text{ca}}(\cdot|q^{L},\mathcal{D}_{q^{L}})),$
$P_{\text{ca}}(d_{i}|q^{L},\mathcal{D}_{q^{L}})=\sum_{h=0}^{H}\sum_{t=0}^{|d_{i}|}\frac{\operatorname{SG}(\operatorname{CA}(0,h,t))}{H}\ |\ d_{i}\in\mathcal{D}_{q^{L}},$
where $\mathcal{D}_{q^{L}}$ is the set of passages returned by the retriever itself and $P_{\text{ca}}$ is the target distribution gathered from the decoder’s cross-attention scores. $\operatorname{SG}$ signifies stop-gradient, which blocks the gradient to ensure the decoder is not affected by the retriever loss, and $\operatorname{CA}$ denotes the cross-attention score at the last decoder layer. The term 0 refers to the first output token, $H$ is the number of cross-attention heads, and $|d_{i}|$ is the length of passage $d_{i}$.
The reader optimises the negative log-likelihood of generating $a^{L}$ given $q^{L}$ and relevant passages $\mathcal{D}_{q^{L}}$ as input: $\mathcal{L}_{\text{reader}}=-P_{\text{ans}}(a^{L}|q^{L},\mathcal{D}_{q^{L}})$. The final loss combines reader and retriever loss: $\mathcal{L}_{\text{e2e}}=\alpha\cdot\mathcal{L}_{\text{KL}}+\mathcal{L}_{\text{reader}}$.
Asynchronous Passage Update.
During training, we need to use the retriever to gather a set of passages $\mathcal{D}_{q^{L}}$ from $\bm{C^{\textbf{Multi}}}$ for each $(q^{L},a^{L})$. [^3] We replace $a^{L}$ with $a_{i}^{En}$ and $\bm{C^{\textbf{Multi}}}$ with $\bm{C^{\textbf{En}}}$ when focusing on cross-lingual retrieval from English corpus. However, since the retriever’s parameters are updated constantly, employing the latest model for retrieval becomes computationally expensive due to the need for recomputing all passage embeddings. To ensure efficient training, we periodically update the retrieved passages for each training query using the most recent model every 1000 steps.